# Mantis: A Versatile Vision-Language-Action Model with Disentangled Visual Foresight

### v1

@@ 4.3 Real World Experiments
We evaluate Mantis on an Agilex platform (Fig. 6 (a)) to investigate the effect of language supervision on preserving the backbone’s capabilities. We design three experimental scenarios, each containing four in-domain (ID) instructions to assess instruction following and four out-of-domain (OOD) instructions to evaluate generalization. Fig. 6 (b) shows the setup of the scenarios and provides example instructions. Effective generalization demands a nuanced understanding and reasoning. For example, the first scenario requires world knowledge (e.g., Taylor Swift), while the second involves basic arithmetic logic. We compare Mantis with $\pi_{0.5}$ [intelligence2025pi_], a state-of-the-art open-source VLA model. Both are fine-tuned separately on the three scenarios, with Mantis maintaining language supervision throughout training. Each instruction is executed 10 times, and performance is reported as the average number of successful executions, allowing up to 5 consecutive attempts per trial.
@@ 5 Conclusion, Limitations and Future Work
Table 3: Comparison on VQA and multimodal understanding benchmarks. Mantis achieves superior performance on 2 of the 3 benchmarks. Compared with the original backbone, its performance decreases only marginally. Bold denotes the best results.
MME OCRBench RealWorldQA
Qwen2.5-VL 2217.3 807 62.1
ECoT 0 12 0
ChatVLA 1435.2 729 57.0
Mantis(Ours) 2070.2 757 56.9
Figure 9: Comparison between Mantis and Mantis-LU.
@@ Appendix B Implementation Details
For visual inputs, video frames and the robot’s primary camera images are cropped and resized to $512\times 512$ pixels, while the robot’s wrist camera images are at $256\times 256$. Both training stage 1 and 2 are trained for one epoch using a cosine learning rate schedule with 500 warm-up steps, a base learning rate of 1e41e-4, and a minimum learning rate of 1e51e-5. For stage 3, a fixed learning rate of 1e51e-5 is used, with vision loss weight $\alpha=0.1$ and language loss weight $\beta=0.005$. Our 38 language-supervision datasets are sourced from those used in the instruction-training stage of LLaVA-OneVision-1.5-Instruct [an2025llava]. We include datasets covering visual question answering, OCR, embodied planning, and other general-purpose tasks, while excluding those focused on chart question answering, medical imaging, and other highly specialized domains. A complete list of the selected datasets is provided in Table 4.
@@ Appendix C Real World Experiments
All in-domain (ID) and out-of-domain (OOD) instructions for the three scenarios are summarized in Table 5. For each scenario, we collected 100 teleoperated demonstration episodes, with 25 episodes per task. Both Mantis and $\pi_{0.5}$ were fine-tuned on the combined dataset containing all four tasks within each scenario. For each model, we adopted the same learning-rate settings used in pretraining stage 1 and trained for 10 epochs. Mantis was supervised with the LLaVA-Instruct dataset [liu2023visual], whereas $\pi_{0.5}$ did not use any language supervision. We set the vision loss weight to $\alpha=0.1$ and the language loss weight for Mantis to $\beta=0.1$. Task-level success counts for both models across the three scenarios are reported in Table 6. Execution examples on real-world tasks is shown in Fig. 10.
@@ Appendix D Language Supervision Ablations
To assess whether language supervision preserves the backbone’s capabilities, we conducted ablation studies comparing Mantis with other VLA models [zhou2025chatvla, zawalski2024robotic] on several VQA and multimodal understanding benchmarks [yin2024survey, liu2024ocrbench]. The original Qwen2.5-VL model was also included for reference, and the results are shown in Table 3. Mantis achieves the best performance on two of the three benchmarks, with only a marginal drop relative to the original backbone, confirming the effectiveness of the language supervision.
@@
Table 4: The 38 datasets used for Mantis language supervision. These datasets are sourced from the instruction-tuning data of LLaVA-OneVision-1.5-Instruct [an2025llava] and cover general visual-language tasks while excluding specialized domains.

### v2

@@ 4.3 Real World Experiments
We evaluate Mantis on an Agilex platform (Fig. 6 (a)) to investigate the effect of language supervision on preserving the backbone’s capabilities. We design three experimental scenarios, each containing four in-domain (ID) instructions to assess instruction following and four out-of-domain (OOD) instructions to evaluate generalization. Fig. 6 (b) shows the setup of the scenarios and provides example instructions. Effective generalization demands a nuanced understanding and reasoning. For example, the first scenario requires world knowledge (e.g., Taylor Swift), while the second involves basic arithmetic logic. We compare Mantis with $\pi_{0.5}$ [21], a state-of-the-art open-source VLA model. Both are fine-tuned separately on the three scenarios, with Mantis maintaining language supervision throughout training. Each instruction is executed 10 times, and performance is reported as the average number of successful executions, allowing up to 5 consecutive attempts per trial.
@@ Appendix B Implementation Details
For visual inputs, video frames and the robot’s primary camera images are cropped and resized to $512\times 512$ pixels, while the robot’s wrist camera images are at $256\times 256$. Both training stage 1 and 2 are trained for one epoch using a cosine learning rate schedule with 500 warm-up steps, a base learning rate of 1e41e-4, and a minimum learning rate of 1e51e-5. For stage 3, a fixed learning rate of 1e51e-5 is used, with vision loss weight $\alpha=0.1$ and language loss weight $\beta=0.005$. Our 38 language-supervision datasets are sourced from those used in the instruction-training stage of LLaVA-OneVision-1.5-Instruct [1]. We include datasets covering visual question answering, OCR, embodied planning, and other general-purpose tasks, while excluding those focused on chart question answering, medical imaging, and other highly specialized domains. A complete list of the selected datasets is provided in Table 4.
@@ Appendix C Real World Experiments
All in-domain (ID) and out-of-domain (OOD) instructions for the three scenarios are summarized in Table 5. For each scenario, we collected 100 teleoperated demonstration episodes, with 25 episodes per task. Both Mantis and $\pi_{0.5}$ were fine-tuned on the combined dataset containing all four tasks within each scenario. For each model, we adopted the same learning-rate settings used in pretraining stage 1 and trained for 10 epochs. Mantis was supervised with the LLaVA-Instruct dataset [29], whereas $\pi_{0.5}$ did not use any language supervision. We set the vision loss weight to $\alpha=0.1$ and the language loss weight for Mantis to $\beta=0.1$. Task-level success counts for both models across the three scenarios are reported in Table 6. Execution examples on real-world tasks is shown in Fig. 10.
@@ Appendix D Language Supervision Ablations
To assess whether language supervision preserves the backbone’s capabilities, we conducted ablation studies comparing Mantis with other VLA models [57, 53] on several VQA and multimodal understanding benchmarks [52, 30]. The original Qwen2.5-VL model was also included for reference, and the results are shown in Table 3. Mantis achieves the best performance on two of the three benchmarks, with only a marginal drop relative to the original backbone, confirming the effectiveness of the language supervision.
@@
Table 4: The 38 datasets used for Mantis language supervision. These datasets are sourced from the instruction-tuning data of LLaVA-OneVision-1.5-Instruct [1] and cover general visual-language tasks while excluding specialized domains.